OpenStudy (ellecullen):

Homework question: Find f(x+h) - f(x) / h for f(x) = -5/x+2 Please help asap

OpenStudy (anonymous):

wtf

OpenStudy (anonymous):

too many numbers im going crazy

OpenStudy (anonymous):

you mean the derivative ? if you mean the derivative , its simply , 0(x+2) + 1 ( -5) / (x+2)^2 -5/(x+2)^2

OpenStudy (anonymous):

0_0

OpenStudy (ellecullen):

I'm not sure what it is called, but the calculation is interesting. Can you walk me through the steps?

OpenStudy (tkhunny):

Homework question: Find f(x+h) - f(x) / h for f(x) = -5/x+2 1) You have not written it correctly. You MEAN [f(x+h) - f(x)] / h The square brackets are NOT optional. 2) You have not written it correctly. You MEAN -5/(x+2) The parentheses are NOT optional. 3) Simply substitute. What's preventing you from doing this? f(x+h) = -5/[(x+h)+2] You already can do that. Now, simplify!

OpenStudy (anonymous):

it is the derivative , instead of putting x , put x+h

OpenStudy (ellecullen):

It sounds complicated, as this question in my homework is in the precalculus topic. I'm having a hard time simplifying.

OpenStudy (tkhunny):

It is NOT the derivative. Do you see a limit in there? @ElleCullen This ratio is very important in early studies of the calculus. This is a surreptitious exposure for you so that it will look familiar when you get there. Seriously, just substitute and simplify. \(\dfrac{f(x+h) - f(x)}{h} = \dfrac{\dfrac{-5}{x+h+2}-\dfrac{-5}{x+2}}{h}\) That's it. Now simplify.

OpenStudy (ellecullen):

let me know if I did the simplification right. f(x+h)-f(x)/h = 0/h/h Is this the way to do it?

OpenStudy (tkhunny):

That makes no sense and you still don't seem to understand the Order of Operations that makes more parentheses necessary. First, write that expression that I gave you. That is where you start. Substitution is all that happened. \(f(x) = \dfrac{-5}{x+2}\) \(f(x+h) = \dfrac{-5}{x+h+2}\)

OpenStudy (ellecullen):

Oh, I see how this goes. I think I can finish this question now :) Elle